### 03 Mar 2026 v1

# Phys4D: Fine-Grained Physics-Consistent 4D Modeling from Video Diffusion

Abstract
Figure 1: Phys4D Our three-stage training paradigm lifts a pretrained video diffusion model into a physics-consistent 4D world model. Bottom-left: a scalable physics-simulation data collection pipeline with diverse camera view control. Top-right: Our 4D physics benchmark and diagnostics that evaluate geometric coherence, motion stability beyond appearance-based metrics.
Recent video diffusion models have achieved impressive capabilities as large-scale generative world models. However, these models often struggle with fine-grained physical consistency, exhibiting physically implausible dynamics over time. In this work, we present Phys4D, a pipeline for learning physics-consistent 4D world representations from video diffusion models. Phys4D adopts a three-stage training paradigm that progressively lifts appearance-driven video diffusion models into physics-consistent 4D world representations. We first bootstrap robust geometry and motion representations through large-scale pseudo-supervised pretraining, establishing a foundation for 4D scene modeling. We then perform physics-grounded supervised fine-tuning using simulation-generated data, enforcing temporally consistent 4D dynamics. Finally, we apply simulation-grounded reinforcement learning to correct residual physical violations that are difficult to capture through explicit supervision. To evaluate fine-grained physical consistency beyond appearance-based metrics, we introduce a set of 4D world consistency evaluation that probe geometric coherence, motion stability, and long-horizon physical plausibility. Experimental results demonstrate that Phys4D substantially improves fine-grained spatiotemporal and physical consistency compared to appearance-driven baselines, while maintaining strong generative performance. Our project page is available at https://sensational-brioche-7657e7.netlify.app/.
Machine Learning, ICML
1 Introduction
Recent video generation models (OpenAI, 2024; Google DeepMind, 2024; Polyak et al., 2024) and current generative world models(Chen et al., 2025c; Yu et al., 2024b, 2023; Huang et al., 2025b) have achieved impressive visual realism and large-scale physics. Yet their success does not imply physical understanding: most are trained to match appearances rather than to model underlying physical state and dynamics. As a result, they frequently fail at fine-grained physics, producing locally inconsistent geometry, unstable object motion, and non-causal dynamics over time (Singer et al., 2022; Zhou et al., 2024). A faithful world model should instead represent how the environment evolves, requiring physics-consistent 4D representations.
A key challenge in physics-consistent 4D world modeling is the lack of scalable supervision for fine-grained physical dynamics. Learning such models requires dense, temporally aligned signals of geometry and motion, which are extremely difficult to obtain from real-world data at scale. Existing 4D datasets(Wen et al., 2025a; Zhai et al., 2024) therefore tend to be appearance-rich but physics-poor, often dominated by camera motion. To address this limitation, we leverage the diversity and controllability of physics-based simulation to obtain large amounts of dynamic video data with accurate ground-truth geometry and motion, providing supervision that is otherwise inaccessible.
However, data alone are insufficient to learn physics-consistent 4D world models. Motivated by the maturity of large-scale video generation pipelines, we ground physical reasoning within models that are both scalable and well pretrained. To this end, we adopt RGB-D video as a 2.5D representation, which explicitly encodes scene geometry and motion while remaining compatible with modern video diffusion architectures. Building on this representation, we introduce Phys4D (Fig. LABEL:fig:head_picturel), which integrates fine-grained physics into video diffusion models through explicit depth supervision for geometric consistency and motion supervision for dynamic coherence. To bring a pretrained video diffusion model to a 4D dynamics model, Phys4D employs a three-stage training pipeline that progressively injects physical structure while preserving generative capacity.
In the first stage, we lift a pretrained video diffusion model from pure 2D appearance modeling toward an explicit 4D spatio-temporal representation by augmenting RGB videos with geometric and motion supervision. Concretely, we apply strong off-the-shelf monocular estimators to annotate large-scale videos generated by the pretrained model, together with carefully curated in-the-wild internet videos. This mixed supervision preserves the robustness of large-scale video generation, while providing dense depth and motion signals that initialize a coherent 4D understanding of scene geometry and dynamics in a domain-agnostic manner. In the second stage, we selectively finetune the high-noise components of the diffusion process with physics-grounded warp-based consistency constraints, coupling geometry and motion to reinforce temporally coherent 3D structure and physically plausible dynamics over long horizons, ultimately transforming the video diffusion model into a physics-consistent 4D world model. However, supervision in the first two stages remains largely pixel- and correspondence-based, which can produce results that are still physically inconsistent at the level of 4D scene dynamics. To address this limitation, the third stage introduces simulation-grounded reinforcement learning, where the model is optimized directly on generated 4D scene trajectories using object-centric, motion-based rewards, explicitly aligning video generation with fine-grained physical outcomes over long horizons.
To evaluate whether a model behaves as a physics-consistent 4D world model, we introduce a comprehensive evaluation protocol that combines content-level physics assessment and 4D world-level physics analysis. Specifically, in addition to established video-based physics benchmarks such as PhysicsIQ(Motamed et al., 2025b), which evaluate physical plausibility from visual and semantic cues, we propose a 4D evaluation suite that probes geometry and motion world representations through trajectory-based analysis. Together, these complementary evaluations assess physical plausibility from what a model shows to how the underlying world evolves, providing a holistic diagnosis of physics-consistent 4D modeling beyond appearance and reconstruction quality.
In summary, our contributions are as follows:
  • We propose Phys4D, a physics-aware training framework for improving fine-grained physical consistency in video diffusion models, focusing on coherent geometry and motion over time.
  • We propose a three-stage training pipeline that integrates physics understanding into video diffusion models.
  • We show how physics-based simulation can be leveraged as a high-fidelity source of geometric, motion, and reward supervision, enabling fine-grained physical alignment that is difficult to obtain from real-world video data alone.
  • We present a set of 4D world consistency diagnostics that evaluate geometric coherence, motion stability, and long-horizon physical plausibility beyond appearance-based video metrics.
2 Related Work
Physics-Aware Video Generation and Evaluation
Recent video generation models(OpenAI, 2024; Google DeepMind, 2024) achieve strong visual realism and temporal coherence, but are largely optimized for appearance rather than physical world modeling. Several works attempt to introduce physical cues into video generation, such as modeling object dynamics(Zhang et al., 2024; Aira et al., 2024) or action-conditioned motion(Chang et al., 2024; Yuan et al., 2025), and a number of benchmarks(Motamed et al., 2025b; Meng et al., 2025; Bansal et al., 2024b) evaluate physical plausibility based on visual or semantic signals. While these approaches can improve or assess appearance-level physical realism, they operate primarily in image space and do not evaluate whether a model maintains coherent geometry, motion, or a consistent physical state over time. In contrast, our work focuses on physics consistency at the level of a 4D world, motivating geometry- and time-aware evaluation beyond appearance-centric video physics assessment. Detailed discussion please refer to Appendix G.1 G.2 G.3
Physics-Aware 4D Modeling
A growing body of work explores incorporating physical priors into generative models, including post hoc physical simulation after generation (Stomakhin et al., 2013b; Chen et al., 2025a; Xie et al., 2024), simulation-guided generation with predefined dynamics (Yuan et al., 2023; Liu and others, 2024a; Li et al., 2025), and learning physical priors from data (Li et al., 2024; Xu et al., 2024; Yang et al., 2025a). These methods typically focus on object-level or short-horizon dynamics, rely on specialized 3D representations, or remain optimization-based, limiting scalability and generalization. In contrast, our work leverages pretrained video diffusion models and treats fine-grained physics as a property of an evolving 4D world. For a detailed discussion, please refer to Appendix G.4
3 Problem Formulation
Given a first-frame RGB image $I_{0}\in\mathbb{R}^{H\times W\times 3}$ and an optional text prompt $p$, our goal is to generate a temporally coherent video together with an explicit, physically consistent 4D world representation that captures the evolving scene geometry and motion over time. Rather than treating video generation as a purely image-level task, we model the underlying scene as a time-varying 3D world whose state evolves across frames.
We represent per-frame scene geometry using depth maps $D_{t}\in\mathbb{R}^{H\times W}$ in the camera coordinate system, and describe scene dynamics through inter-frame motion. Specifically, we model 2D motion using optical flow $F_{t\rightarrow t+1}\in\mathbb{R}^{H\times W\times 2}$, and, when necessary, consider full 3D motion in the form of scene flow $S_{t\rightarrow t+1}\in\mathbb{R}^{H\times W\times 3}$, which captures per-pixel 3D displacement in camera space.
Our model learns a conditional generative mapping
$(I_{0},p)\ \longmapsto\ \Big(\{I_{t}\}_{t=0}^{T},\ \{D_{t}\}_{t=0}^{T},\ \{F_{t\rightarrow t+1}\}_{t=0}^{T-1}\Big),$
optionally producing $\{S_{t\rightarrow t+1}\}_{t=0}^{T-1}$, where the generated video frames are treated as observations of an underlying 4D scene evolution. Beyond visual realism, we require world-level physical consistency, such that geometry and motion predictions remain mutually consistent over time (e.g., warping $D_{t}$ under $F_{t\rightarrow t+1}$ agrees with $D_{t+1}$), and collectively describe a physically plausible evolution of the scene state. More detailed discussion is given at App  F
4 Simulation Data Collection
We present a physics-based simulation pipeline designed to address the limited physical fidelity and diversity of existing training data for 4D world modeling. In contrast to prior datasets that focus on narrow physical regimes or specific simulation paradigms (Huang et al., 2025a; Chang et al., 2024), Phys4D supports large-scale data generation with diverse object types, rich physical interactions, and accurate geometry and motion supervision. Figure 2 provides an overview of the simulation pipeline and the range of supported object categories.
Figure 2: Overview of Phys4D simulation framework. (a)Left A Large-scale async data collection pipeline. (b)Right Supported physical objects: rigid and articulated structures, garments, fluids, thermodynamics, deformable, inflatables, ropes, and granular materials.
Diverse Object Types and Physics Models
Based on Isaac Sim (NVIDIA Corporation, 2024; Lu et al., 2024), our simulation pipeline supports a broad range of object categories and physical phenomena to enable fine-grained physical interactions. We adopt tailored simulation methods according to object-specific physical characteristics: rigid and articulated objects are modeled with rigid-body dynamics, large garments and fluid materials are simulated using Position-Based Dynamics  (Bender et al., 2015), and small elastic garments (e.g., gloves and socks) as well as deformable everyday objects (e.g., toys and sponges) are simulated using the Finite Element Method  (Ciarlet, 2002). Additional object types such as ropes, inflatables, granular materials, and thermodynamic effects are handled by appropriate physics solvers, collectively exposing the model to diverse fine-grained physical regimes. Detailed simulation method please refer to Appendix H
Systematic Physical Parameterization
To disentangle physical phenomena for effective learning, we decompose generation into nine distinct categories (e.g., Rigid Body, Fluids, Soft Bodies, Thermodynamics). This supports curriculum learning where models progress from mastering individual laws to complex compositional scenarios. As detailed in Table 1, we enforce systematic randomization across material, environmental, and geometric attributes. We structure the datasets with progressive complexity (single to multi-object) (Bengio et al., 2009) and introduce parameter perturbation to prevent homogeneity, resulting in over 20000 unique physical configurations.
Table 1: Simulation Parameter & Curriculum Configuration. We structure the dataset with progressive complexity and randomize physical attributes to ensure diverse and robust data generation.
Category Parameters Range / Configuration
Curriculum Complexity Dist. Single (30%), Two-body (35%), Multi-object (35%)
Material Density ($\rho$), Young’s Mod. ($E$) $\rho\in[100,10\,000]\text{ kg/m}^{3}$, $E\in[10^{6},10^{11}]\text{ Pa}$
Poisson ($\nu$), Friction ($\mu_{s,k}$) $\nu\in[0.2,0.5]$, $\mu_{s,k}\in[0.1,1.0]$
Restitution ($e$) $e\in[0.1,0.95]$
Environment Gravity ($g$), Lighting $g\in[5,15]\text{ m/s}^{2}$, Intensity $\in[100,100\,000]\text{ lux}$
Variables Air viscosity, Temperature gradients, 1000+ HDRIs
Geometry Shape Types Cubes, spheres, cylinders, complex meshes
Scale ($s$), Aspect Ratio $s\in[0.1,5.0]\text{ m}$, Randomized aspect ratios
Perturbation Perturbation Ratio ($\delta$) $\delta\in[0.05,0.25]$ (Randomized around base values)
Visuals Reflectance (Roughness, Metallic) Roughness $\in[0.05,0.95]$, Metallic $\in[0,1]$
Appearance Randomized RGB $\in[0,1]^{3}$, Surface textures
Multi-Modal Camera Strategies
We employ two complementary camera strategies. Fixed multi-view cameras provide dense spatial coverage by placing multiple static cameras around each scene, while dynamic trajectory cameras simulate realistic cinematography by sampling continuous camera motions along parameterized trajectories, producing frame-level pose annotations.
Asynchronous Parallel Data Collection
To overcome the single-scene bottleneck of standard simulators, we implement asynchronous parallel execution. By eliminating global synchronization, multiple independent scenes run concurrently within a unified instance, allowing immediate resource recycling upon scene completion. This architecture achieves near-linear scaling efficiency.
[…]
5 Method
5.1 Method Overview
Phys4D converts a pretrained video diffusion model into a physics-consistent 4D world model via a three-stage training paradigm (Fig. 3). Starting from a standard text- and image-conditioned video diffusion architecture, we attach lightweight auxiliary heads for depth and motion prediction, enabling explicit supervision of scene geometry and dynamics. The three stages—pretraining (Sec 5.2), supervised fine-tuning (Sec 5.3), and reinforcement learning (Sec 5.4) —progressively enforce geometry–motion consistency and fine-grained physical plausibility.
Figure 3: Overview of the Phys4D training pipeline. Our three-stage paradigm progressively injects physics into a pretrained video diffusion model. Stage 1 (blue): The DiT backbone is frozen while depth and motion heads are trained on pseudo-labeled RGB videos. Stage 2 (green): The backbone is adapted via LoRA using physics simulation data with ground-truth annotations; a warp consistency loss couples depth and motion predictions. Stage 3 (orange): Generated RGB-D-motion outputs are lifted to 4D point clouds and compared against simulator ground-truth via 4D Chamfer Distance; PPO optimization uses this reward to correct residual physical violations. The output is a generated video with temporally consistent depth maps and motion fields.
5.2 Pseudo-Supervised Pretraining
Data Sources.
In the first stage, we bootstrap geometry and motion representations using large-scale pseudo-supervision. We employ off-the-shelf depth (Chen et al., 2025b) and optical flow (Teed and Deng, 2020) estimators to annotate (i) videos generated by the pretrained video diffusion model and (ii) curated internet videos, preserving visual diversity while injecting geometric and motion cues. Detailed data info please refer to Appendix A.2
Model Architecture. We build on a pretrained video diffusion model with a DiT-style backbone and augment it with two lightweight auxiliary heads for explicit geometry and motion prediction. A motion head is attached to diffusion features via a forward hook at the DiT output layer, consisting of LayerNorm followed by a linear projection, and predicts a patchified motion representation that is unpatchified into a multi-channel optical flow field for adjacent frames. In parallel, a depth head with the same feature interface predicts per-frame depth maps $D_{t}\in\mathbb{R}^{H\times W}$; following (Chen et al., 2025b), the depth head incorporates temporal attention and multi-scale feature fusion to encourage temporally coherent depth estimation across video frames. During this stage, the DiT backbone is frozen and only the auxiliary heads are optimized. For Details, please refer to Supp B.1
Training Objective.
We adopt flow matching to parameterize the underlying video generative process. Flow-based generative models learn a time-dependent velocity field $v_{\theta}(\mathbf{z}_{t},t)$ that transforms noise samples $\mathbf{z}_{0}\sim\mathcal{N}(0,I)$ into data samples $\mathbf{z}_{1}\sim q$ by solving the ordinary differential equation
$\frac{d\mathbf{z}_{t}}{dt}=v_{\theta}(\mathbf{z}_{t},t).$ (1)
Following the flow matching framework (Lipman et al., 2023; Liu et al., 2023), the model is trained by directly regressing the predicted velocity field onto a target conditional vector field $u_{t}(\cdot\mid\mathbf{z}_{1})$:
$\mathcal{L}_{\mathrm{FM}}=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{1}}\Big[\|v_{\theta}(\mathbf{z}_{t},t)-u_{t}(\mathbf{z}_{t}\mid\mathbf{z}_{1})\|_{2}^{2}\Big],$ (2)
where $\mathbf{z}_{t}$ is sampled from a predefined conditional probability path. A commonly used choice is linear interpolation between noise and data:
$\mathbf{z}_{t}=t\mathbf{z}_{1}+(1-t)\mathbf{z}_{0},\qquad u_{t}(\mathbf{z}_{t}\mid\mathbf{z}_{1})=\mathbf{z}_{1}-\mathbf{z}_{0}.$ (3)
To enable explicit supervision over scene geometry and motion, we attach auxiliary depth and motion heads to the frozen diffusion backbone. Given diffusion features at time $t$, the heads predict per-frame depth $\hat{D}_{t}$ and inter-frame optical flow $\hat{F}_{t\rightarrow t+1}$. The supervision signals $D_{t}$ and $F_{t\rightarrow t+1}$ are obtained as described above. The auxiliary heads are trained using supervised regression losses:
$\mathcal{L}_{\text{depth}}=\sum_{t}\ell(D_{t},\hat{D}_{t}),\\ \mathcal{L}_{\text{motion}}=\sum_{t}\ell(F_{t\to t+1},\hat{F}_{t\to t+1}).$ (4)
where $\ell$ denotes a combination of L1 and $\hat{D}_{t}$, $\hat{F}_{t\to t+1}$ are ground-truth depth and optical flow respectively.
During Stage I, the video backbone is kept fixed and only the auxiliary heads are optimized, yielding a stable initialization for subsequent physics-aware fine-tuning stages. Detailed prelinminary and proof please refer to Supp F.1
5.3 Physics-Grounded Supervised Fine-Tuning
In the second stage, we perform physics-grounded supervised fine-tuning using simulation-generated data with accurate ground-truth geometry and motion. Unlike pseudo labels from real videos, simulation provides physically valid supervision derived from underlying dynamics, enabling direct enforcement of world-level consistency. We selectively fine-tune the high-noise regime of the diffusion model using lightweight LoRA adapters, while continuing to train the auxiliary heads.
Overall Objective.
The Stage II objective augments the original diffusion loss with physics-aware consistency terms:
$\mathcal{L}=\mathcal{L}_{\text{FM}}+\lambda_{w}\mathcal{L}_{\text{warp}}+\lambda_{\text{rgb}}\mathcal{L}_{\text{rgb-warp}},$ (5)
where $\mathcal{L}_{\text{FM}}$ denotes the rectified flow matching loss used by the base video generation model (Eq 2). Supervised depth and motion losses (Eq 4) remain active as defined in Stage I, but are omitted here for brevity.
Warp-Based Geometry–Motion Consistency.
Per-frame supervision alone does not guarantee temporal coherence of the underlying 3D structure. To explicitly couple geometry and motion across time, we introduce a warp-based consistency loss:
$\mathcal{L}_{\text{warp}}=\sum_{t}\big\|\mathcal{W}(D_{t},F_{t\rightarrow t+1})-D_{t+1}\big\|_{1},$ (6)
where $\mathcal{W}(\cdot)$ denotes backward warping using the predicted motion field. This loss enforces that the depth map at time $t$, when transported according to the predicted motion, agrees with the depth at time t+1t{+}1, directly operationalizing temporal geometry–motion consistency. We optionally include an RGB warp loss $\mathcal{L}_{\text{rgb-warp}}$ to encourage photometric consistency via flow-based warping.
By enforcing physics-grounded temporal consistency using simulation supervision, this stage transforms the model from frame-wise estimation toward a coherent 4D representation.
5.4 Simulation Grounded RL Fine-Tuning
The final stage applies reinforcement learning to correct 4D residual physical inconsistencies that persist after supervised fine-tuning and are difficult to capture through 2D differentiable losses. While Stage II enforces geometry–motion consistency at the pixel level, subtle violations of object motion and interaction dynamics may still arise. To address this gap, we leverage simulation-provided visual ground truth to define reward signals that explicitly align generated videos with physically valid motion patterns.
Denoising as Sequential Decision Making.
Following prior work (Wallace et al., 2023; Liu et al., 2025), we view the reverse denoising process as a sequential decision process, where each denoising step corresponds to an action that transitions the latent state toward the final video sample. For details please refer to the Supp  F.2
Flow-SDE for Stochastic Exploration.
The flow-matching video model in Eq. (1) defines a deterministic probability flow ODE, which produces a single sampling trajectory and thus lacks stochastic exploration. To introduce exploration while preserving the same marginals, we adopt the probability flow formulation (Song et al., 2020) and convert the ODE into an equivalent SDE:
$d\mathbf{z}_{t}=\left(v_{\theta}(\mathbf{z}_{t},t)-\frac{1}{2}g^{2}(t)\,\nabla\log q_{t}(\mathbf{z}_{t})\right)dt+g(t)\,d\mathbf{w}_{t},$ (7)
where $q_{t}(\mathbf{z}_{t})$ denotes the marginal density and $\mathbf{w}_{t}$ is a standard Wiener process.
Leveraging the rectified-flow score–velocity identity, we substitute the induced closed-form score into Eq. (7) and choose $g(t)=\sigma_{t}$, yielding the final Flow-SDE:
$d\mathbf{z}_{t}=\left[v_{\theta}(\mathbf{z}_{t},t)+\frac{\sigma_{t}^{2}}{2t}\left(\mathbf{z}_{t}+(1-t)\,v_{\theta}(\mathbf{z}_{t},t)\right)\right]dt+\sigma_{t}\,d\mathbf{w}_{t}.$ (8)
Discretizing Eq. (8) induces a Gaussian transition $p_{\theta}(\mathbf{z}_{t-\Delta t}\mid\mathbf{z}_{t})$, which we interpret as a stochastic denoising policy for RL fine-tuning. Detailed derivations are provided in Supp. F.3.
Reward from Simulation-Grounded 4D Motion Consistency.
To encourage fine-grained physical plausibility beyond image-level alignment, we define a terminal reward that directly operates on the generated 4D scene dynamics. Given a generated video, we first lift them into a sequence of spatio-temporal point clouds using the predicted geometry and motion. This yields a set of generated 4D points $\mathcal{P}^{\text{gen}}=\{(x,y,z,\tau)\}$ describing object trajectories over time. In parallel, we obtain a physically ground-truth 4D point set $\mathcal{P}^{\text{phys}}$ from simulation. We then define the reward as the negative 4D Chamfer Distance between the two point sets:
$r(\mathbf{z}_{0},c)=-\mathrm{CD}_{4\mathrm{D}}\!\left(\mathcal{P}^{\text{gen}},\,\mathcal{P}^{\text{phys}}\right),$ (9)
where the 4D distance jointly measures spatial deviation and temporal misalignment. This reward explicitly penalizes physically implausible object trajectories and aligns generation with coherent 4D scene evolution. Detailed Implementation please refer to Supp B.2
Policy Optimization.
We optimize the resulting stochastic policy using Proximal Policy Optimization (PPO). For brevity, we omit the optimization objective and training details here and provide a full description in the Supp  F.4
Figure 4: Qualitative comparison on Physics-IQ scenarios. We compare Wan2.2-5b baseline (middle) with Wan2.2-5b + Phys4D (right) across three physical interaction types: object placement on a rotating platform (top), ball rolling dynamics (middle), and fluid pouring (bottom). Phys4D produces more consistent object geometry, physically plausible motion, and stable temporal dynamics compared to the baseline, which exhibits shape distortion and incoherent physical behavior.
6 4D World Model Benchmark
Table 2: Performance on the Physics-IQ benchmark. We report reconstruction error (MSE), motion consistency metrics (IoU), and the overall Physics-IQ score. Commercial models only report the aggregate score.
Group Model Params MSE $\downarrow$ ST-IoU $\uparrow$ S-IoU $\uparrow$ WS-IoU $\uparrow$ Score (%) $\uparrow$
Ours CogVideoX + Phys4D 5B 0.009 0.169 0.252 0.157 30.2
WAN2.2 + Phys4D 5B 0.014 0.107 0.214 0.122 25.6
Open-Sora V1.2 + Phys4D 1.1B 0.016 0.098 0.195 0.112 22.4
Open-source CogVideoX 5B 0.013 0.116 0.222 0.142 18.8
WAN2.2 5B 0.016 0.088 0.150 0.105 16.8
Open-Sora V1.2 1.1B 0.021 0.072 0.135 0.092 14.5
Commercial VideoPoet / Pika 1.0 / Sora 20.3 / 13.0 / 10.0
In order to comprehensively evaluate 4D world model generation quality, our benchmark is organized along two axes: (i) video-based physics assessment, which measures physical plausibility from visual and semantic cues using established benchmarks, and (ii) 4D world-level evaluation, which examines the consistency of geometry, motion, and temporal evolution in the generated scene.
6.1 Video-Based Physics Evaluation
Benchmark.
We evaluate on Physics-IQ (Motamed et al., 2025a), a benchmark designed to assess physical understanding in generative video models. The benchmark comprises 198 test scenarios covering diverse physical phenomena. Each scenario includes three camera viewpoints and two natural variations, yielding 1,188 test videos.
Metrics.
Following the standard Physics-IQ protocol, we evaluate physical consistency through four complementary measures: SpatialIoU metrics and MSE, which quantifies pixel-level reconstruction error. Lower MSE and higher IoU values indicate better physical plausibility.
Implementation.
We apply Phys4D to three backbone: WAN2.2-5b (Wan et al., 2025), CogVideoX-5b (Yang et al., 2025b), and Open-Sora V1.2 (Zheng et al., 2024). For evaluation, we generate 5-second videos conditioned on the switch frame following the Physics-IQ protocol.
Results.
Fig. 4 and Table 2 presents our results on the Physics-IQ benchmark. Phys4D consistently improves the overall Physics-IQ score across multiple backbones. On CogVideoX-5b, Phys4D increases the score from 18.8 to 30.2 (+11.4 absolute), with similarly strong gains on WAN2.2-5b (16.8 $\rightarrow$ 25.6) and Open-Sora V1.2 (14.5 $\rightarrow$ 22.4). These results demonstrate that our physics-grounded training paradigm robustly enhances physical understanding in an architecture-agnostic manner, substantially elevating open-source video diffusion models.
Table 3: Per-frame 3D Geometry and Video Quality. Our model demonstrates superior geometric correctness (lower AbsRel) while maintaining high visual quality (lower FVD). DAV2 denotes DepthAnythingV2.(Yang et al., 2024)
Depth Accuracy Video Quality
Model AbsRel $\downarrow$ RMSE $\downarrow$ $\delta_{1}$ $\uparrow$ $\delta_{2}$ $\uparrow$ $\delta_{3}$ $\uparrow$ FVD $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
WAN + DAV2 0.3929 0.9012 26.00 52.69 75.95 140.02 84.02 20.80
CogVideoX + DAV2 0.3483 0.7854 32.45 68.97 84.35 130.03 84.98 20.34
Phys4D + DAV2 0.3694 0.8429 27.38 55.42 80.29 135.46 84.93 20.76
Phys4D (Ours) 0.2711 0.6604 43.64 80.04 95.77 124.52 85.98 21.89
6.2 4D World Consistency Evaluation
Figure 5: Qualitative Results on 4D Experiment. From left to right: the ground-truth 4D point cloud and the generated 4D point cloud at $1/4$ of the sequence; the generated 4D point cloud at $1/2$ of the sequence; the novel-time generated 4D point cloud at $3/4$ of the sequence; and the generated 4D point cloud at the final frame. This visualization highlights the model’s ability to maintain coherent geometry and object motion over long horizons, as well as to interpolate consistent 4D structure at unseen timestamps.
Table 4: Full 4D World Evaluation. We report 4D Chamfer Distance (Geometry), Worldline/Trajectory metrics (Dynamics), and Novel Time Interpolation error (Continuity). Detailed metrics info please refer to E.3
Geometry Trajectory / Worldline Dynamics Novel Time
Model Chamfer $\downarrow$ L2 Error $\downarrow$ Mean Drift $\downarrow$ Final Drift $\downarrow$ Fail Rate $\downarrow$ Length $\uparrow$ Depth Err $\downarrow$ Warp Err $\downarrow$
WAN + OTF 0.5058 0.5367 0.5369 0.5576 12.38% 83.18 0.5841 1.1076
CogVideoX + OTF 0.4923 0.5239 0.5239 0.5348 11.34% 84.67 0.5432 1.1254
OpenSora + OTF 0.5286 0.5694 0.5637 0.5834 13.54% 80.24 0.6534 1.1573
Phys4D (Ours) 0.4626 0.4928 0.4829 0.5078 9.54% 87.21 0.5340 1.1288
To assess whether Phys4D internalizes a coherent world representation beyond pixel-level synthesis, we introduce a three-tier 4D world-level evaluation protocol that probes spatiotemporal reasoning at increasing temporal scales: per-frame 3D geometry, short-term geometry and motion consistency, and long-horizon 4D world evolution. This hierarchy disentangles frame-level, local temporal, and global world-level errors, enabling a comprehensive evaluation of physically plausible 4D scene consistency over time.
Per-frame 3D Geometry Accuracy.
A coherent 4D world model must first capture accurate instantaneous 3D geometry. We evaluate per-frame depth accuracy by comparing predicted depth maps against simulation ground truth under a fixed camera setting. This isolates geometric understanding from temporal reasoning. As shown in Table 3, Phys4D achieves better depth accuracy across all metrics while maintaining strong video quality. These results indicate that Phys4D learns more accurate per-frame 3D structure without sacrificing perceptual fidelity.
Temporal Geometry & Motion Consistency.
Beyond per-frame geometry, a physical world model must maintain spatiotemporal self-consistency between geometry, appearance, and motion. We evaluate this property using two complementary warp-based tests: (i) depth warping with ground-truth flow and (ii) RGB/latent warping with predicted flow. Table 5 summarizes depth warp errors, RGB warp reconstruction, and optical flow accuracy. Phys4D consistently achieves lower warp errors and substantially better motion metrics, indicating improved alignment between geometry and motion over time. These results show that Phys4D produces temporally coherent dynamics rather than frame-wise plausible but inconsistent motion.
World-level Evolution As shown in Table 4, We evaluate global 4D geometry using the 4D Chamfer Distance. For object dynamics, we construct worldlines by first sampling pixels at the initial frame, tracking their 2D trajectories over time using ground-truth optical flow, and back-projecting these trajectories into 3D using the predicted depth maps. The resulting 3D trajectories are compared against ground-truth worldlines obtained from simulation, yielding worldline and trajectory errors. We further test temporal continuity at unseen timestamps. Given only observed frames, the model reconstructs a partial 4D scene and is queried at novel time steps via flow-based interpolation. We measure novel-time depth and warp errors to evaluate whether the inferred geometry remains consistent across time, indicating a continuous underlying $(x,y,z,t)$ world rather than frame-local predictions. Overall, Phys4D yields more stable worldlines and lower trajectory drift while maintaining accurate global geometry. Detailed Experiment setting please refer to Appendix E.3. Qualitative result is shown on Fig. 5
Table 5: Temporal Warp and Motion Consistency. We report reconstruction errors for both depth and RGB warping, alongside optical flow endpoint error (EPE). OTF refers to the optical flow estimation method from SEA-RAFT  (Wang et al., 2024)
Temporal Warp Consistency Motion Dynamics
Depth Warp RGB Warp
Model L1 $\downarrow$ Charb. $\downarrow$ LPIPS $\downarrow$ Charb. $\downarrow$ EPE $\downarrow$ Fl-all $\downarrow$ 1px Out $\downarrow$
WAN+OTF 0.7990 0.7898 0.2409 0.0507 1.2516 6.2173 15.69
CogVideoX+OTF 0.7054 0.7893 0.2564 0.0527 1.2343 4.3243 12.34
OpenSora+OTF 0.7436 0.8534 0.2783 0.0589 1.2718 7.2343 18.96
Phys4D (Ours) 0.5141 0.5240 0.1818 0.0518 0.5054 1.3013 4.35
6.3 Ablation Studies
We ablate key components of Phys4D to understand their individual contributions to physical consistency. Table 6 reports an ablation study on the effect of each training stage measured by Physics-IQ. Starting from the pretrained baseline, Stage 2 yields the largest performance gain by incorporating simulation-grounded supervision and warp-based consistency. Stage 3 further improves physical plausibility through reinforcement learning. Overall, the monotonic improvements across stages validate the effectiveness of our curriculum-style training pipeline. More ablation studies please refer to Appendix D.2
Table 6: Ablation on training stages. Each stage contributes incrementally to physical consistency, with Stage 2 (physics-grounded fine-tuning) providing the largest improvement.
Configuration Score (%) $\Delta$
Base (WAN2.2-5b) 16.8
+ Stage 2 (Physics-Grounded SFT) 22.6 +5.8
+ Stage 3 (RL Fine-tuning) 25.6 +3.0
7 Conclusion
We presented Phys4D, a three-stage framework that lifts pretrained video diffusion models into physics-consistent 4D world models. Our results highlight the importance of explicit 4D world-level training and evaluation for advancing generative world models.
8 Impact Statements
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix Overview
This appendix provides supplementary materials that support the main paper and improve clarity, completeness, and reproducibility. Specifically, it expands on data construction, extended experiments, qualitative analyses, theoretical details, and physics-based simulation settings that are not fully described in the main text due to space constraints.
Figure 6: Qualitative Result Of Phys4D Generated Scene.
Appendix A: Data.
We describe the data sources used throughout training, including internet videos, model-generated samples, and simulation-generated data. This section details data preprocessing, filtering strategies, and annotation procedures for depth and motion supervision, clarifying how diverse supervision signals are constructed across different stages of training.
Appendix B: Implementation
This appendix provides implementation details for the proposed Phys4D framework, covering data preparation, model training, and reinforcement learning fine-tuning. Our implementation follows the three-stage training paradigm described in the main paper and is designed to progressively lift video diffusion models from appearance-level generation to physically consistent 4D world modeling.
Appendix C: Qualitative Results.
We present additional qualitative results to complement the quantitative evaluations in the main paper. This section includes visual demonstrations of generated videos, depth maps, and motion fields, highlighting improvements in fine-grained physical consistency, temporal coherence, and long-horizon stability.
Appendix D: Experiments.
We report extended experimental results and ablation studies that are omitted from the main paper for brevity. These experiments further analyze the effects of different training stages, supervision signals, and design choices, providing deeper insight into the behavior of the proposed method.
Appendix E: 4D Physics Consistency Benchmarks and Metrics.
We provide additional details on benchmark settings and evaluation protocols. This section elaborates on the computation of 4D world consistency metrics and discusses their relationship to standard appearance-based video evaluation measures.
Appendix F: Preliminaries and Theory.
We include mathematical preliminaries and theoretical details underlying our approach, including flow matching, deterministic ODE sampling, and the stochastic formulations used for reinforcement learning fine-tuning. This section contains derivations and clarifications that support the methodology presented in the main paper.
Appendix G: Additional Related Work.
We discuss related work that could not be fully covered in the main paper, providing further context on physics-aware video generation, 4D modeling, and learning-based world representations.
Appendix H: Physics Simulation.
We describe the physics simulation settings used for supervision and reward construction, including scene configurations, object properties, and motion extraction procedures. This section explains how simulation is leveraged as a source of high-fidelity geometric and motion information for fine-grained physical alignment.
Appendix A Data
A.1 Simulation Data
To enable large-scale 4D data collection at high throughput and low cost, we build our simulation pipeline on an asynchronous parallel execution framework. Our system collects data from a large number of simulation sub-environments concurrently, significantly improving GPU utilization and simulation throughput compared to sequential or synchronously parallel simulators. This design allows us to generate large-scale, diverse simulation data at substantially higher speed and lower computational cost.
Tiled Rendering for Efficient Visual Simulation.
To further improve rendering efficiency, we adopt tiled camera rendering, where multiple camera views from different simulation sub-environments are rendered within a single large framebuffer. This approach amortizes rendering overhead across environments and avoids redundant GPU context switching. We follow the tiled rendering strategy provided by Isaac Lab (NVIDIA Isaac Lab, ), which enables scalable visual data generation with minimal performance degradation as the number of parallel environments increases.
Asynchronous Execution.
While parallelism is critical for simulator efficiency, naïve synchronous execution leads to significant idle time when sub-environments exhibit heterogeneous dynamics. In vision-based control and data collection, a policy typically performs inference once to generate a target pose or waypoint, after which the robot executes a trajectory of variable length. Synchronizing such trajectories across environments results in unnecessary blocking and under-utilization of computational resources.
To address this challenge, our simulator supports fully asynchronous execution. Each policy-driven motion is encapsulated as an atomic action with a well-defined interface, allowing high-level policies to invoke and monitor actions independently across environments. This design eliminates global synchronization barriers and maximizes hardware utilization, enabling efficient closed-loop data collection under heterogeneous execution.
Data Management and Asynchronous Control.
Supporting asynchronous execution is non-trivial, as the underlying simulator operates in a SIMD (Single Instruction, Multiple Data) paradigm, whereas asynchronous execution requires MIMD (Multiple Instruction, Multiple Data) semantics. To bridge this gap, we introduce a motion-planning-based abstraction layer that translates high-level asynchronous commands into low-level, time-continuous trajectories. This abstraction allows each sub-environment to progress independently while preserving simulator stability.
Scene Management and Heterogeneous Environments.
To support heterogeneous parallel simulation, different sub-environments may require distinct scene configurations. We address this by introducing a Layout Manager that programmatically controls asset placement across environments. The Layout Manager imports user-specified objects with configurable randomization, enabling diverse, controllable, and reproducible scene layouts. This design allows us to efficiently scale data collection across a wide range of scene configurations within a single simulator instance.
Multi-Modal Camera Strategies
We employ two complementary camera strategies to balance static scene coverage and dynamic viewpoint variation. Fixed multi-view cameras provide dense spatial coverage by placing multiple static cameras around each scene, capturing aligned RGB images together with calibrated intrinsics and extrinsics and dense geometric annotations such as depth and optical flow. In parallel, dynamic trajectory cameras simulate realistic cinematography by sampling continuous camera motions (e.g., orbit, dolly, crane, tracking, and first-person) along parameterized trajectories, producing frame-level pose annotations for training trajectory-aware video generation models.
Simulation Throughput and Data Scale.
Our data collection pipeline is optimized for high-throughput simulation. On a single NVIDIA RTX 4090 GPU with 24 GB memory, we run up to 1,024 simulation environments in parallel. Each environment is simulated for 10 seconds of effective interaction time, resulting in approximately 5000 video clips collected per hour on a single GPU. This efficiency enables large-scale data generation at low computational cost.
We construct a base set of 200 simulation scenes and expand it to approximately 250,000 distinct environments through domain randomization over initial object configurations, physical parameters, and surface textures. For each environment, we collect five camera trajectories, including three fixed-view cameras and two dynamic camera motions, providing diverse spatial and temporal observations of the same underlying physical scene.
A.2 Internet Video Pretraining Data
Motivation.
To improve the robustness and generalization of the depth and motion heads, we leverage large-scale real-world videos for pseudo-supervised pretraining. Compared to simulation data, Internet videos exhibit substantially greater diversity in scene layout, appearance, and motion patterns, which is critical for learning domain-agnostic geometry and motion representations. This pretraining stage equips the model with strong visual and geometric priors before introducing physics-grounded supervision.
Data Sources.
We assemble the pretraining corpus using a hybrid collection strategy that combines automated web crawling with selective manual curation. Videos are sourced from multiple public platforms, including YouTube, Pexels, Pixabay, Mixkit, and Bilibili, ensuring broad coverage of real-world environments and camera motions.
Automated Quality Filtering.
To maintain data quality at scale, we apply a multi-stage automated filtering pipeline. We first use the LAION aesthetic scorer (Schuhmann and others, 2022) to remove visually low-quality or noisy clips. Next, we employ an OCR-based text detector based on PaddleOCR (Liao and others, 2022) to identify and discard videos containing prominent overlaid text, watermarks, or subtitles that may interfere with depth and motion estimation. Finally, videos that are extremely short, corrupted, or visually inconsistent are automatically removed. This filtering process yields a large-scale, diverse, and high-quality video dataset suitable for robust geometry and motion pretraining.
Final Dataset Curation.
Starting from an initial pool of approximately 100,000 candidate video clips, our automated filtering pipeline retains about 5,000 high-quality videos that satisfy our fine-grained requirements. These selected clips exhibit sufficient motion diversity, clear object boundaries, and stable visual quality, making them suitable for robust depth and motion pretraining. This curated subset forms the final Internet video dataset used in our pseudo-supervised pretraining stage.
Appendix B Implementation
B.1 Model Architecture
We augment the pretrained DiT backbone with two auxiliary prediction heads that extract physics-aware signals from the diffusion features.
Motion Head. We attach a motion prediction head to DiT features captured via a forward hook at the DiT output layer. The head applies LayerNorm followed by a linear projection with modulation, producing patchified motion outputs that are unpatchified to a 4-channel motion field $(dx,dy,c_{2},c_{3})$ per frame, where $(dx,dy)$ represents optical flow and $(c_{2},c_{3})$ encode auxiliary motion information.
Depth Head. We attach a depth prediction head following the spatiotemporal head design of Video Depth Anything (Chen et al., 2025b). The head builds on the DPT architecture with four temporal attention layers inserted at low-resolution stages to enable cross-frame reasoning. Each temporal layer applies multi-head self-attention along the temporal dimension with absolute positional embeddings. The head uniformly samples four feature maps from the DiT backbone, processes them through Reassemble and Fusion layers to produce a multi-scale feature pyramid, and outputs per-frame depth maps $D_{t}\in\mathbb{R}^{H\times W}$.
B.2 RL Reward Implementation: Simulation-Grounded 4D Chamfer
4D point distance and 4D Chamfer.
We measure spatio-temporal alignment between two 4D points $p=(x,y,z,\tau)$ and $q=(x^{\prime},y^{\prime},z^{\prime},\tau^{\prime})$ in the camera coordinate system using
$d_{4\mathrm{D}}(p,q)=\left\|(x,y,z)-(x^{\prime},y^{\prime},z^{\prime})\right\|_{2}^{2}+\alpha^{2}\,(\tau-\tau^{\prime})^{2},$ (10)
where $\alpha$ converts a unit time difference (one frame) into an equivalent spatial scale (meters). Given two 4D point sets $\mathcal{P}^{\text{gen}}$ and $\mathcal{P}^{\text{gt}}$, the symmetric 4D Chamfer Distance is defined as
$\mathrm{CD}_{4\mathrm{D}}(\mathcal{P}^{\text{gen}},\mathcal{P}^{\text{gt}})=\frac{1}{|\mathcal{P}^{\text{gen}}|}\sum_{p\in\mathcal{P}^{\text{gen}}}\ \min_{q\in\mathcal{P}^{\text{gt}}}d_{4\mathrm{D}}(p,q)+\frac{1}{|\mathcal{P}^{\text{gt}}|}\sum_{q\in\mathcal{P}^{\text{gt}}}\ \min_{p\in\mathcal{P}^{\text{gen}}}d_{4\mathrm{D}}(p,q).$ (11)
We define the terminal RL reward as the negative 4D Chamfer Distance:
$r(\mathbf{z}_{0},c)=-\mathrm{CD}_{4\mathrm{D}}\!\left(\mathcal{P}^{\text{gen}},\mathcal{P}^{\text{gt}}\right).$ (12)
Generated 4D point construction.
The model outputs video frames $\{I_{t}\}_{t=0}^{T-1}$, depth maps $\{D_{t}\}_{t=0}^{T-1}$ (in camera coordinates), and motion predictions in the form of 2D optical flow $\{F_{t\rightarrow t+1}\}_{t=0}^{T-2}$ or 3D scene flow $\{S_{t\rightarrow t+1}\}_{t=0}^{T-2}$. We construct the generated 4D point set $\mathcal{P}^{\text{gen}}$ as follows.
(i) Depth unprojection. Using the camera intrinsics $K$, each pixel $(u,v)$ at frame $t$ is back-projected to 3D:
$\mathbf{X}^{\text{gen}}_{t}(u,v)=D_{t}(u,v)\,K^{-1}\!\begin{bmatrix}u\\ v\\ 1\end{bmatrix}\in\mathbb{R}^{3}.$ (13)
We form a 4D point by appending the discrete timestamp $\tau_{t}\in\{0,\dots,T-1\}$: $p^{\text{gen}}_{t}(u,v)=(\mathbf{X}^{\text{gen}}_{t}(u,v),\tau_{t})$.
(ii) Moving-object selection via motion. To focus the reward on object dynamics, we select moving regions using the predicted motion. When using optical flow, we compute a magnitude map from $F_{t\rightarrow t+1}$ and define
$M^{\text{gen}}_{t}(u,v)=\mathbb{I}\big(\|F_{t\rightarrow t+1}(u,v)\|_{2}>\delta\big).$
When 3D scene flow $S_{t\rightarrow t+1}$ is available, we analogously threshold its magnitude. Only 4D points with $M^{\text{gen}}_{t}(u,v)=1$ are retained, forming $\mathcal{P}^{\text{gen}}$ over all frames.
Simulation-derived ground-truth 4D points.
From simulation, we obtain ground-truth 3D point clouds and motion in the world coordinate system. For each frame $t$, ground-truth points $\mathbf{X}^{\text{gt,w}}_{t}$ are transformed into the camera coordinate system using camera extrinsics $(R_{t},\mathbf{t}_{t})$:
$\mathbf{X}^{\text{gt}}_{t}=R_{t}\,\mathbf{X}^{\text{gt,w}}_{t}+\mathbf{t}_{t}.$ (14)
We construct 4D points $p^{\text{gt}}_{t}=(\mathbf{X}^{\text{gt}}_{t},\tau_{t})$ with the same discrete timestamp $\tau_{t}$. Moving-object points are selected using simulator-provided ground-truth motion (e.g., per-point 3D displacement), thresholded by the same $\delta$, yielding the set $\mathcal{P}^{\text{gt}}$.
Temporal weight $\alpha$.
The temporal weight $\alpha$ balances spatial error (meters) and temporal misalignment (frames) in Eq. (10). As our environments are room-scale, we set $\alpha=0.03\,\text{m}$, corresponding to the typical per-frame displacement of moving objects in the simulator. This choice makes a one-frame temporal mismatch comparable to a few centimeters of spatial error, and we observe stable behavior under moderate variations of $\alpha$.
Appendix C Qualitative Result
Figure 7 and Figure 4 qualitatively compare Phys4D with the base video diffusion model across a diverse set of physical interaction scenarios under static camera settings, including rigid-body manipulation, multi-object dynamics, soft-body deformation, fluid flow, and contact-rich interactions. While the baseline model often generates visually coherent frames, it frequently violates object-level physical constraints, leading to shape distortion, object duplication, incorrect interaction outcomes, or implausible motion patterns over time.
In manipulation scenarios, such as object placement and grasping, baseline models tend to introduce non-physical deformations or geometry drift, even for nominally rigid objects. Phys4D, in contrast, preserves object integrity throughout the interaction, maintaining consistent geometry and contact behavior as objects are manipulated or released. For multi-object scenes, baseline generations commonly suffer from object count inconsistencies and unstable trajectories, whereas Phys4D maintains correct object identities and produces smooth, coherent motion consistent with the underlying scene layout.
We further observe substantial improvements in fluid–object interactions. In pouring scenarios, baseline models often violate basic physical principles, such as gravity alignment or container geometry, resulting in fluid emerging from incorrect locations. Phys4D generates fluid motion that is directionally consistent with gravity and constrained by the container structure, yielding more realistic flow behavior over time.
Across all scenarios, these qualitative results highlight that Phys4D goes beyond frame-wise visual plausibility, enforcing object-level continuity, interaction consistency, and physically grounded dynamics. This supports the claim that Phys4D captures a coherent 4D world representation, rather than synthesizing isolated, appearance-consistent frames.
Figure 7: Qualitative Result Of Phys4D Generated Scene.
Appendix D Experiment
D.1 Physics-IQ
For evaluation, we generate 5-second videos conditioned on the switch frame following the Physics-IQ protocol. We use nucleus sampling with temperature 1.0 and top-$p$ = 0.95 for all experiments. Phys4D achieves substantial improvements over base video diffusion models across multiple backbones. When applied to CogVideoX-5b, Phys4D improves the Spatiotemporal IoU from 0.116 to 0.169 (+45.7% relative) and the Spatial IoU from 0.222 to 0.252 (+13.5% relative), while simultaneously reducing MSE from 0.013 to 0.009 (30.8% reduction). Similarly, Phys4D boosts WAN2.2-5b’s Spatiotemporal IoU from 0.088 to 0.107 (+21.6% relative) and Open-Sora V1.2 from 0.072 to 0.098 (+36.1% relative). These consistent improvements across different backbones demonstrate that our physics-grounded training paradigm is architecture-agnostic and effectively transfers to various video diffusion models.
Notably, Phys4D (CogVideoX-5b) achieves the best performance across all metrics, with a Spatiotemporal IoU of 0.169, Spatial IoU of 0.252, and Weighted Spatial IoU of 0.157, substantially outperforming all baseline models. These results indicate that our simulation-grounded training approach can elevate open-source models to achieve significantly improved physical understanding.
D.2 Ablation
Loss Components. Table 7 ablates individual loss terms in our training objective. The warp consistency loss $\mathcal{L}_{\text{warp}}$ provides the most significant contribution by explicitly coupling geometry and motion predictions, enforcing that warped depth maps align temporally. Removing depth supervision ($\mathcal{L}_{\text{depth}}$) or motion supervision ($\mathcal{L}_{\text{motion}}$) both degrade performance substantially, confirming that joint supervision over both signals is necessary for physics-consistent generation. The RGB warp loss ($\mathcal{L}_{\text{rgb-warp}}$) provides complementary photometric consistency, with its removal causing moderate performance degradation.
Table 7: Ablation on loss components. Warp consistency and joint depth-motion supervision are critical for physics-consistent generation.
Configuration Physics-IQ (%)
Full model 25.6
   w/o $\mathcal{L}_{\text{warp}}$ 21.3
   w/o $\mathcal{L}_{\text{depth}}$ 19.8
   w/o $\mathcal{L}_{\text{motion}}$ 20.5
   w/o $\mathcal{L}_{\text{rgb-warp}}$ 24.1
Appendix E Benchmark
We introduce a structured 4D benchmark Fig. 6 to evaluate whether a video generation model internalizes a coherent $(x,y,z,t)$ world representation beyond per-frame appearance. The benchmark is organized into three complementary evaluation tiers that progressively increase the level of spatiotemporal reasoning required.
The first tier assesses per-frame 3D geometry accuracy, measuring whether each generated frame exhibits correct static geometry under a fixed camera setting. The second tier evaluates temporal geometry and motion consistency, probing whether predicted geometry remains self-consistent across adjacent frames under motion. The final tier focuses on explicit 4D reconstruction, testing the model’s ability to recover continuous 4D scene structure and object trajectories over extended time horizons, including reasoning at unseen timestamps.
Together, these three tiers provide a comprehensive and diagnostic evaluation of geometry, motion, and world-level coherence, enabling fine-grained analysis of where and how video generation models succeed or fail in representing dynamic 4D worlds.
E.1 Per-frame 3D Geometry Evaluation
This experiment evaluates whether the generated videos exhibit accurate per-frame 3D geometry under a fixed camera setting. Our goal is to assess static geometric correctness independently of temporal effects, focusing on whether each frame reflects proper 3D scene understanding rather than purely appearance-level realism.
E.1.1 Experimental Setting
We evaluate per-frame geometry using simulation data with ground-truth depth annotations. For each method, we generate videos conditioned on the same initial frames and prompts, and extract the predicted depth maps from the corresponding depth heads. All evaluations are conducted frame-wise, without enforcing temporal consistency, to isolate static 3D geometry accuracy.
To ensure a fair comparison, all models are evaluated on identical test scenes with fixed camera intrinsics and viewpoints. Video quality metrics are computed on the generated RGB frames, while depth accuracy is measured by comparing predicted depth against simulation ground truth.
E.1.2 Evaluation Metrics
We assess per-frame 3D geometry accuracy using standard depth estimation metrics, including Absolute Relative Error (AbsRel), Root Mean Squared Error (RMSE), and threshold accuracy metrics $\delta_{1}$, $\delta_{2}$, and $\delta_{3}$. Lower AbsRel and RMSE indicate more accurate depth prediction, while higher $\delta$ scores reflect better geometric correctness.
In addition to geometric accuracy, we report common video quality metrics to verify that improvements in depth estimation do not come at the cost of visual fidelity. Specifically, we measure Fréchet Video Distance (FVD), Structural Similarity Index (SSIM), and Peak Signal-to-Noise Ratio (PSNR) on the generated RGB videos.
E.1.3 Results and Analysis
Table 3 reports per-frame depth accuracy and video quality results. Phys4D achieves substantially lower AbsRel and RMSE compared to all baselines, indicating more accurate per-frame 3D geometry reconstruction. Notably, Phys4D also attains significantly higher $\delta_{1}$, $\delta_{2}$, and $\delta_{3}$ scores, suggesting improved geometric correctness across a wide range of depth thresholds.
Importantly, these gains in geometric accuracy do not compromise visual quality. Phys4D maintains competitive or superior FVD, SSIM, and PSNR scores relative to baseline video generation models. These results demonstrate that Phys4D improves per-frame 3D understanding while preserving high-quality video synthesis, validating the effectiveness of our geometry-aware training strategy.
E.2 Temporal Geometry & Motion Consistency
This experiment evaluates whether the predicted geometry and motion remain temporally self-consistent across adjacent frames. Under a fixed camera setting, temporal consistency provides a strong test of whether the model captures coherent scene evolution rather than producing individually plausible but temporally inconsistent frames.
E.2.1 Experimental Setting
We evaluate temporal geometry and motion consistency on video sequences generated by different models. For baseline methods, depth and optical flow are obtained using off-the-shelf estimators (OTF), specifically DepthAnythingV2 for depth and RAFT for optical flow. Phys4D directly predicts depth and motion using its learned geometry and motion heads.
All evaluations are conducted under fixed camera intrinsics and viewpoints. Temporal consistency is assessed between consecutive frames by warping predicted quantities from time $t$ to t+1t{+}1 and comparing them against corresponding targets. This setting isolates temporal coherence from viewpoint changes and emphasizes consistency in geometry and motion prediction.
E.2.2 Evaluation Metrics
We consider two complementary aspects of temporal consistency.
Depth Warp Consistency. We warp the predicted depth map $D_{t}^{\mathrm{pred}}$ to the next frame using ground-truth optical flow $F_{t\rightarrow t+1}^{\mathrm{gt}}$, and compare it against the ground-truth depth $D_{t+1}^{\mathrm{gt}}$. Warping error is measured using both L1 and Charbonnier losses. To ensure a rigorous evaluation, errors are reported separately for occluded and non-occluded regions.
RGB / Latent Warp Consistency. We evaluate appearance-level temporal coherence by warping RGB frames or latent features using the predicted optical flow $F_{t\rightarrow t+1}^{\mathrm{pred}}$. The warped result is compared against the next-frame target using LPIPS and Charbonnier distances. Latent-space metrics are included to provide a more stable measure of temporal consistency that is less sensitive to pixel-level noise.
E.2.3 Results and Analysis
Table 5 summarizes temporal warp consistency and motion dynamics results. Phys4D achieves substantially lower depth warp errors than all baselines, indicating more consistent geometric prediction across frames. This improvement holds for both L1 and Charbonnier losses and across occluded and non-occluded regions.
Phys4D also outperforms baseline methods in RGB and latent warp consistency, achieving lower LPIPS and Charbonnier distances. These results suggest that Phys4D produces smoother and more coherent temporal evolution at both geometric and appearance levels.
In addition, Phys4D exhibits significantly improved motion dynamics, with lower endpoint error (EPE), reduced outlier rates, and fewer large-motion failures. Overall, these results demonstrate that Phys4D enforces strong temporal self-consistency in geometry and motion, beyond what can be achieved by post-hoc off-the-shelf depth and flow estimation.
E.3 Full 4D Reconstruction Evaluation
This experiment evaluates whether a model represents an explicit 4D world with coherent geometry and dynamics, rather than a sequence of independently plausible frames. Given generated videos, we lift predicted depth and motion into a unified $(x,y,z,t)$ representation and assess geometric accuracy, object-level dynamics, and temporal continuity over extended horizons.
E.3.1 Experimental Setting
All evaluations are conducted under a fixed-camera setting. Using predicted depth and motion, each video is lifted into a spatio-temporal point cloud or trajectory representation in camera coordinates. Simulation ground truth provides aligned depth, optical flow, and scene flow, allowing direct comparison in 3D space and time. Baselines rely on off-the-shelf depth and flow estimators (OTF), while Phys4D directly predicts geometry and motion through learned heads.
E.3.2 4D Geometry Consistency
Overview.
This sub-experiment evaluates whether the model reconstructs a globally consistent 4D scene geometry across space and time.
Metric.
We measure 4D Chamfer Distance between predicted and ground-truth spatio-temporal point sets. Each 4D point is represented as $(x,y,z,\tau)$, where $\tau$ denotes normalized time. The metric penalizes both spatial misalignment and temporal inconsistency, providing a holistic measure of 4D geometric accuracy.
Results and Analysis.
As shown in Table 4, Phys4D achieves the lowest 4D Chamfer Distance among all methods, indicating more accurate and temporally aligned 4D geometry. This suggests that Phys4D maintains coherent scene structure over time, rather than reconstructing geometry independently at each frame.
E.3.3 Worldline and Trajectory Dynamics
Overview.
This sub-experiment evaluates object-level motion consistency by examining 3D worldlines over time. The goal is to test whether a model preserves coherent object trajectories in $(x,y,z,t)$ space, rather than producing temporally unstable geometry that appears plausible only at individual frames.
Worldline Construction.
We adopt a reviewer-safe protocol to construct worldlines that is independent of the model’s predicted motion. At the initial frame $t=1$, we uniformly sample $N$ pixels $\{p_{1,i}\}_{i=1}^{N}$ from the valid object mask or depth-valid region. Each pixel is then tracked in 2D using ground-truth optical flow:
$p_{t+1,i}=p_{t,i}+F^{\mathrm{gt}}_{t\rightarrow t+1}(p_{t,i}),$ (15)
yielding a stable and unambiguous 2D trajectory $\{p_{t,i}\}_{t=1}^{T}$.
The 2D trajectories are lifted into 3D using the model-predicted depth:
$\mathbf{x}^{\mathrm{pred}}_{t,i}=\Pi^{-1}(p_{t,i},D^{\mathrm{pred}}_{t}),$ (16)
where $\Pi^{-1}$ denotes back-projection using known camera intrinsics. Ground-truth 3D worldlines $\mathbf{x}^{\mathrm{gt}}_{t,i}$ are constructed analogously using ground-truth depth or scene flow from simulation.
Metrics.
We report several complementary worldline metrics. The Worldline L2 Error measures the average 3D deviation over time:
$\mathrm{WorldlineErr}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{T}\sum_{t=1}^{T}\left\|\mathbf{x}^{\mathrm{pred}}_{t,i}-\mathbf{x}^{\mathrm{gt}}_{t,i}\right\|_{2}.$ (17)
To characterize temporal accumulation of errors, we report the Trajectory Drift as a function of time:
$\mathrm{Drift}(t)=\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{x}^{\mathrm{pred}}_{t,i}-\mathbf{x}^{\mathrm{gt}}_{t,i}\right\|_{2}.$ (18)
We summarize this curve using both the mean drift and the final-time drift.
We additionally report a Failure Rate, defined as the fraction of worldlines whose deviation exceeds a threshold $\tau$ at any time step, indicating track breakage or physically implausible motion. Finally, we report the average valid trajectory length, which reflects how long worldlines remain stable before failure.
We report several complementary worldline metrics: Worldline L2 Error measures the average 3D positional deviation between predicted and ground-truth trajectories across all time steps. Mean Drift quantifies the average accumulated deviation over time, while Final Drift captures long-horizon error at the end of the sequence. Fail Rate measures the fraction of trajectories that deviate beyond a predefined threshold, indicating track breakage or physically implausible motion. Finally, Trajectory Length reports the average valid duration of tracked worldlines, reflecting temporal stability.
Results and Analysis.
As shown in Table 4, Phys4D achieves lower worldline error, reduced drift accumulation, and fewer trajectory failures than all baselines. These results indicate improved long-horizon motion coherence and object identity preservation in 3D space.
E.3.4 Novel-Time Interpolation
Overview.
This sub-experiment evaluates temporal continuity by testing whether the learned representation supports coherent geometry at unseen timestamps. Unlike C.3, which evaluates consistency at observed frames, C.4 probes whether the model encodes a continuous underlying 4D world.
Protocol.
Given a video sequence, we split the timeline into observed and novel timestamps:
$\mathcal{T}_{\mathrm{obs}}=\{0,2,4,\ldots\},\qquad\mathcal{T}_{\mathrm{novel}}=\{1,3,5,\ldots\}.$ (19)
Only observed frames are used to construct the 4D representation. For each novel time $t+1\in\mathcal{T}_{\mathrm{novel}}$, we estimate geometry via flow-based interpolation.
Specifically, we warp predicted depth from the previous observed frame:
$\tilde{D}_{t+1}=\mathrm{warp}\!\left(D^{\mathrm{pred}}_{t},\,F^{\mathrm{pred}}_{t\rightarrow t+1}\right),$ (20)
where $F^{\mathrm{pred}}_{t\rightarrow t+1}$ is the predicted optical flow. This yields an interpolated depth estimate at the novel timestamp.
Metrics.
We report two primary metrics. The Novel-Time Depth Error measures geometric accuracy at unseen times:
$\mathrm{DepthErr}_{\mathrm{novel}}=\frac{1}{|\mathcal{T}_{\mathrm{novel}}|}\sum_{t\in\mathcal{T}_{\mathrm{novel}}}\left\|\tilde{D}_{t}-D^{\mathrm{gt}}_{t}\right\|_{1}.$ (21)
The Novel-Time Warp Error evaluates temporal consistency by comparing the interpolated depth against a motion-warped reference:
$\mathrm{WarpErr}_{\mathrm{novel}}=\left\|\mathrm{warp}(D^{\mathrm{pred}}_{t-1},F^{\mathrm{gt}}_{t-1\rightarrow t})-\tilde{D}_{t}\right\|_{1}.$ (22)
Lower values indicate better temporal continuity of the reconstructed 4D scene.
Results and Analysis.
Phys4D achieves lower novel-time depth error than all baselines (Table 4), indicating more accurate interpolation at unseen timestamps. Warp error remains comparable to strong baselines, suggesting that Phys4D maintains temporal consistency while improving geometric fidelity.
Appendix F Preliminary and Theory
F.1 Flow Matching for Video Generation
Flow-based generative models provide a continuous alternative to diffusion models for learning complex data distributions. Similar to diffusion models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020), flow generative models aim to transform samples from a simple noise distribution into data samples by learning a time-dependent velocity field that governs a deterministic probability flow (Papamakarios et al., 2021; Song and others, 2021; Xu et al., 2022; Lipman et al., 2023; Liu et al., 2023). Specifically, given an initial noise sample $\mathbf{z}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and a target data sample $\mathbf{z}_{1}\sim q(\mathbf{z})$, flow-based models define a continuous trajectory $\mathbf{z}_{t}$ evolving according to an ordinary differential equation (ODE):
$\frac{d\mathbf{z}_{t}}{dt}=v_{t}(\mathbf{z}_{t}),$ (23)
where $v_{t}$ denotes a time-dependent velocity field parameterized by a neural network.
Recent works propose flow matching as a simulation-free training framework for learning such velocity fields (Albergo and Vanden-Eijnden, 2023). Instead of solving stochastic differential equations or estimating score functions, flow matching directly regresses the model-predicted velocity field onto a target conditional vector field that defines a probability path toward data samples. Following the flow matching framework (Lipman et al., 2023; Liu et al., 2023), the model is trained by minimizing
$\mathcal{L}_{\mathrm{FM}}=\mathbb{E}_{t,\mathbf{z}_{0},\mathbf{z}_{1}}\Big[\|v_{\theta}(\mathbf{z}_{t},t)-u_{t}(\mathbf{z}_{t}\mid\mathbf{z}_{1})\|_{2}^{2}\Big],$ (24)
where $u_{t}(\cdot\mid\mathbf{z}_{1})$ is a predefined conditional vector field and $\mathbf{z}_{t}$ is sampled from the corresponding conditional probability path.
A commonly used and effective choice is linear interpolation between noise and data:
$\mathbf{z}_{t}=t\mathbf{z}_{1}+(1-t)\mathbf{z}_{0},\qquad u_{t}(\mathbf{z}_{t}\mid\mathbf{z}_{1})=\mathbf{z}_{1}-\mathbf{z}_{0},$ (25)
which induces a Gaussian conditional distribution $\mathbf{z}_{t}\sim\mathcal{N}(t\mathbf{z}_{1},(1-t)^{2}\mathbf{I})$. Notably, flow matching is not restricted to Gaussian endpoints or linear paths and can be generalized to interpolate between arbitrary distributions (Albergo and Vanden-Eijnden, 2023; Liu et al., 2023). This flexibility has recently enabled efficient large-scale generative modeling, including applications to high-dimensional and structured data such as videos, where flow-based formulations offer favorable trade-offs between sampling efficiency and temporal coherence (Xu et al., 2022; Lin and others, 2024b; Peng and others, 2025).
F.2 Denoising as a Markov Decision Process
We view the iterative denoising procedure of video generative models as a sequential decision-making problem. Following prior work on reinforcement learning for generative diffusion and flow models (Wallace et al., 2023; Ren et al., 2024), we cast the reverse generative process as a finite-horizon Markov Decision Process (MDP).
At each timestep, the state is defined as
$s_{t}\triangleq(c,t,\mathbf{z}_{t}),$ (26)
where $c$ denotes the conditioning prompt, $t$ the current reverse-time index, and $\mathbf{z}_{t}$ the latent variable. An action corresponds to producing the next latent state,
$a_{t}\triangleq\mathbf{z}_{t-1},$ (27)
and the policy is induced by the generative model,
$\pi_{\theta}(a_{t}\mid s_{t})\triangleq p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},c).$ (28)
The transition dynamics are deterministic and enforce consistency between consecutive states:
$P(s_{t-1}\mid s_{t},a_{t})=\delta(c)\,\delta(t-1)\,\delta(\mathbf{z}_{t-1}),$ (29)
where $\delta(\cdot)$ denotes the Dirac delta distribution. Trajectories terminate after $T$ denoising steps, starting from the initial distribution
$\rho_{0}(s_{T})=p(c)\,\delta(T)\,\mathcal{N}(\mathbf{0},\mathbf{I}).$ (30)
Rewards are only assigned upon completion of the denoising process:
$R(s_{t},a_{t})=\begin{cases}r(\mathbf{z}_{0},c),&t=0,\\ 0,&\text{otherwise},\end{cases}$ (31)
where $r(\mathbf{z}_{0},c)$ is given by our simulation grounded reward function. This formulation enables direct optimization of sample-level objectives without modifying the generative architecture.
F.3 Stochasticizing Flow-Based Generative Processes
Flow matching and rectified flow models define deterministic sampling trajectories governed by ordinary differential equations (ODEs). Specifically, the reverse-time evolution of the latent variable is given by
$d\mathbf{z}_{\tau}=v_{\theta}(\mathbf{z}_{\tau},\tau)\,d\tau,$ (32)
where $v_{\theta}$ is the learned velocity field. While such deterministic dynamics yield efficient sampling, they are insufficient for policy exploration in reinforcement learning.
To enable exploration while preserving the marginal distribution of generated samples, we convert the ODE into an equivalent stochastic differential equation (SDE). This construction follows the probability flow principle, which establishes that a suitable SDE can share the same marginal densities as the original deterministic flow (Song and others, 2021; Liu et al., 2023; Albergo and Vanden-Eijnden, 2023).
The resulting Flow-SDE takes the form
$d\mathbf{z}_{\tau}=\left(v_{\theta}(\mathbf{z}_{\tau},\tau)-\frac{1}{2}g^{2}(\tau)\,\nabla\log q_{\tau}(\mathbf{z}_{\tau})\right)d\tau+g(\tau)\,d\mathbf{w}_{\tau},$ (33)
where $q_{\tau}$ denotes the marginal density of $\mathbf{z}_{\tau}$, $g(\tau)$ controls the noise magnitude, and $\mathbf{w}_{\tau}$ is a Wiener process.
For rectified flows, the score function admits a closed-form expression in terms of the velocity field (Liu et al., 2023):
$\nabla\log q_{\tau}(\mathbf{z}_{\tau})=-\frac{1}{\tau}\big(\mathbf{z}_{\tau}+(1-\tau)v_{\theta}(\mathbf{z}_{\tau},\tau)\big).$ (34)
Substituting this relation into Eq. (7) and choosing $g(\tau)=\sigma_{\tau}$, we obtain
$d\mathbf{z}_{\tau}=\Big[v_{\theta}(\mathbf{z}_{\tau},\tau)+\frac{\sigma_{\tau}^{2}}{2\tau}\big(\mathbf{z}_{\tau}+(1-\tau)v_{\theta}(\mathbf{z}_{\tau},\tau)\big)\Big]d\tau+\sigma_{\tau}\,d\mathbf{w}_{\tau}.$ (35)
This stochastic formulation introduces controlled exploration into the generative process while retaining the original marginal distributions, making it well suited for reinforcement learning fine-tuning.
F.4 Policy Optimization with PPO
Given the MDP formulation and the stochastic sampling dynamics, we optimize the generative policy using Proximal Policy Optimization (PPO) (Schulman and others, 2017). At each timestep, the policy ratio is defined as
$r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}\mid s_{t})}{\pi_{\theta_{\mathrm{old}}}(a_{t}\mid s_{t})},$ (36)
and the clipped PPO objective is
$\mathcal{L}_{\mathrm{PPO}}=\mathbb{E}_{t}\Big[\min\big(r_{t}(\theta)\hat{A}_{t},\mathrm{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\big)\Big],$ (37)
where $\hat{A}_{t}$ denotes the estimated advantage. This objective stabilizes policy updates and has been shown effective for fine-tuning diffusion- and flow-based generative models (Black and others, 2023; Ren and others, 2024; Liu and others, 2025).
Appendix G Related Work
G.1 Large-Scale Video Generation and World Models
Recent large-scale video generation models, such as Sora (OpenAI, 2024), Veo (Google DeepMind, 2024), and MovieDiffusion (Polyak et al., 2024), have achieved impressive visual realism and long-horizon coherence, motivating their use as foundations for generative world modeling. Recent world model systems further extend these capabilities toward persistent and interactive environments, either by explicitly constructing 3D scenes or by augmenting video generation with memory and feedback mechanisms (Chen et al., 2025d; Agarwal and others, 2025; Yu et al., 2024a; Wang and others, 2025; Team and others, 2025). While these approaches emphasize scalability, global coherence, and interactivity, the underlying physical state is often implicit, making fine-grained physical consistency—such as local geometry–motion coupling and precise object interactions—difficult to enforce. In contrast, our work explicitly targets fine-grained physics as a core objective, complementing large-scale world modeling with geometry- and motion-consistent 4D representations.
G.2 4D Datasets for Dynamic World Modeling
Datasets that jointly capture spatial geometry and temporal dynamics are essential for studying 4D world modeling. Early efforts focus on controlled environments or specific interaction settings. For example, HOI4D provides large-scale egocentric RGB-D sequences with dynamic point clouds, semantic labels, and pose annotations for category-level human–object interaction analysis (Liu et al., 2022). Similarly, BEHAVE and InterCap capture human–object interactions using multi-view RGB-D systems, enabling detailed reconstruction of articulated motion and object dynamics (Bertiche et al., 2022; Huang et al., 2022). While these datasets offer high-quality annotations, they are limited in scene diversity and scale.
To improve realism and coverage, several datasets adopt multi-view capture systems to reconstruct dynamic scenes in a 4D setting. Panoptic Studio and subsequent extensions provide synchronized multi-camera recordings for full-body motion capture and free-viewpoint reconstruction (Joo et al., 2015). More recent datasets such as Neu3D and Dynamic3DScenes leverage multi-view video to recover time-varying geometry and appearance of dynamic scenes (Lin et al., 2023; Li et al., 2023). These datasets support 4D reconstruction benchmarks but are typically restricted to studio environments and short temporal spans.
Recent works further scale up 4D datasets by exploiting in-the-wild videos and automatic annotation pipelines. DynamicVerse introduces a large-scale multimodal 4D dataset derived from internet videos, providing metric-scale geometry, camera parameters, instance-level masks, motion cues, and hierarchical textual descriptions across over 100K dynamic scenes (Wen et al., 2025b). Complementary datasets such as OmniWorld and Ego-Exo4D extend this direction by covering diverse environments, viewpoints, and sensor modalities, facilitating evaluation of general-purpose 4D world understanding models (Zhou et al., 2025; Grauman et al., 2023). Collectively, these datasets form the foundation for benchmarking models that jointly reason about geometry, motion, and semantics in dynamic real-world scenes.
G.3 Benchmarks for Physics-Aware Video Generation
To facilitate standardized evaluation, several benchmarks have been proposed to assess general video generation quality. VBench (Huang and others, 2024) and EvalCrafter (Liu and others, 2024b) provide comprehensive evaluation frameworks covering dimensions such as visual fidelity, diversity, temporal coherence, and scalability. T2V-CompBench (Sun and others, 2024) further focuses on compositionality, testing whether generated videos faithfully follow complex and multi-object textual prompts. While these benchmarks are effective for measuring appearance-level and semantic alignment, they do not explicitly evaluate physical plausibility.
Recent studies highlight the importance of physical realism in video generation and propose dedicated physics-oriented benchmarks. VideoPhy (Bansal et al., 2024a) evaluates physical commonsense reasoning in generated videos, while PhyGenBench (Meng and others, 2024) assesses adherence to basic physical laws such as object permanence and gravity. DEVIL (Liao and others, 2024) emphasizes dynamic realism, focusing on motion continuity and temporal consistency, and Physics-IQ (Motamed et al., 2025b) further probes understanding of fundamental physical principles including fluid dynamics and thermodynamics. Despite these advances, existing benchmarks are often limited in scenario diversity or focus on a narrow set of physical laws.
Motivated by these limitations, recent efforts introduce substantially expanded benchmarks that evaluate video generation models across a broader range of physical phenomena, including both realistic and deliberately unphysical dynamics (Meng and others, 2025). Empirical results show that state-of-the-art models perform significantly worse under identical settings, revealing the increased difficulty and diagnostic value of physics-aware benchmarks and underscoring the need for models that reason about physical consistency beyond visual realism.
G.4 Physics-Aware Generation
To address the lack of physical plausibility in video generation, recent works have explored incorporating explicit physical knowledge into generative pipelines. Depending on when and how physical priors are injected, existing methods can be broadly categorized into three types.
Generation then Physical Simulation.
This line of work first generates static images or 3D assets using generative models, and subsequently applies external physics simulators to animate them into dynamic scenes or videos. Representative approaches employ particle- or continuum-based simulators, such as the Material Point Method (MPM), to model object deformation and motion (Stomakhin et al., 2013a). Recent studies adopt this paradigm to produce physically plausible dynamics by post-processing generated content with physics engines (Lin and others, 2024a; Hsu and others, 2024). While this approach offers explicit and controllable physical behavior, it typically requires substantial manual effort and task-specific simulator design, limiting scalability and generalization.
Physical Simulation then Generation.
Another category of methods first performs physical simulation to obtain plausible motion trajectories or object states, which are then used as conditional inputs for video generation models. For example, PhysGen (Lin and others, 2024a) segments dynamic objects from an input image, simulates their motion using Newtonian mechanics, and conditions a video generator on both simulated trajectories and static backgrounds. Although this strategy improves controllability and motion realism, the generative model itself does not possess intrinsic physical reasoning capability, and physical rules and parameters must be manually specified for each scenario.
Generation with Learned Physics Priors.
More recent approaches aim to incorporate physics priors directly into the generative process by leveraging learned or pretrained models (Chang et al., 2024; Zhang and others, 2025; Chefer and others, 2025). These methods encode physical knowledge as latent dynamics, auxiliary supervision, or physics-aware evaluators. For instance, PhysT2V (Xue et al., 2025) employs large language models and vision–language models to iteratively evaluate and refine generated videos for improved physical consistency. However, such approaches often rely on implicit physical understanding acquired through data fitting, which may fail under physically challenging or out-of-distribution scenarios. Our method broadly follows this paradigm while grounding learned physics priors with physics-clean supervision to improve robustness and generalization.
Appendix H Physics Simulation
H.1 Simulation Methodology
Our simulation pipeline incorporates multiple physics solvers to model diverse physical phenomena encountered in complex interactive scenes, including cloth, fluids, deformable objects, rigid bodies, and environmental flow. Each component is selected to balance physical realism, numerical stability, and large-scale data generation efficiency.
H.1.1 Cloth Simulation via Position-Based Dynamics
We simulate garments using a Position-Based Dynamics (PBD) formulation, which represents cloth as a set of particles connected by geometric constraints. The cloth surface is discretized into a triangular mesh, where each vertex corresponds to a particle with associated position, velocity, and mass properties. Physical behavior emerges from enforcing constraints that regulate stretching, bending, and collisions.
Stretching constraints preserve local distances between neighboring particles, preventing excessive elongation, while bending constraints maintain angular relationships across adjacent mesh elements to model resistance to folding. Collision constraints ensure non-penetration between cloth particles and other objects in the scene. Rather than integrating forces explicitly, PBD iteratively projects particle positions to satisfy these constraints at each time step, yielding stable and efficient cloth dynamics even under large deformations.
H.1.2 Fluid Simulation with Particle-Based Constraints
Fluid behavior is simulated using a particle-based PBD framework, where each particle represents a small volume of fluid. The primary objective is to enforce approximate incompressibility while maintaining stable interactions with solid boundaries.
This is achieved by imposing density constraints that regulate local particle concentration, together with collision constraints that handle fluid–solid interactions. At each simulation step, particles are advanced under external forces, then iteratively corrected to satisfy density and collision constraints. This approach enables visually plausible fluid motion with high numerical stability, making it suitable for large-scale data generation.
H.1.3 Deformable Object Simulation via Finite Elements
For volumetric deformable objects, we employ a Finite Element Method (FEM) formulation that captures elastic and damping effects under external forces. Objects are discretized into tetrahedral meshes, with material behavior governed by elasticity, density, and damping parameters.
The simulation separates meshes used for visualization, collision detection, and physical computation, allowing efficient resolution of contact and deformation. By solving the equations of motion over finite elements, the simulator produces realistic volumetric deformations while maintaining controllable computational cost.
H.1.4 Environmental Flow and Wind Effects
To model environmental forces such as wind, we incorporate flow-based models that approximate the interaction between moving air and scene objects. These models capture directional flow, turbulence, and pressure-induced forces that influence both rigid and deformable bodies.
Wind effects are specified through configurable flow fields that define direction and magnitude, and their influence is applied consistently across interacting objects. This enables dynamic responses such as cloth fluttering and object deflection under varying environmental conditions.
H.1.5 Rigid Body Dynamics
Rigid objects are simulated using standard rigid body dynamics, where objects are assumed to maintain fixed internal geometry. Their motion is governed by linear and angular dynamics under external forces and collision impulses.
Collision handling relies on geometric approximations such as convex hulls or signed distance representations to balance accuracy and efficiency. Constraint solvers ensure stable resolution of contacts and joint limits, allowing reliable simulation of complex multi-object interactions.
H.2 Parameterization and Physical Diversity
Rather than tuning individual parameters for specific outcomes, we adopt broad parameter ranges to encourage physical diversity across scenes. Within each physics module, parameters controlling stiffness, damping, friction, density, and interaction strength are randomized within physically plausible bounds.
This strategy enables the simulator to generate a wide spectrum of behaviors for cloth, fluids, deformable objects, and rigid bodies, while avoiding overfitting to narrow physical regimes. As a result, the collected data exhibits substantial variability in geometry, motion patterns, and interaction dynamics, which is critical for training robust 4D world models.

### 06 Mar 2026 v2

@@
 1 Introduction
-Recent video generation models (OpenAI, 2024; Google DeepMind, 2024; Polyak et al., 2024) and current generative world models(Chen et al., 2025c; Yu et al., 2024b, 2023; Huang et al., 2025b) have achieved impressive visual realism and large-scale physics.
+Recent video generation models (OpenAI, 2024; Google DeepMind, 2024; Polyak et al., 2024) and current generative world models(Chen et al., 2025c; Yu et al., 2024, 2023; Huang et al., 2025b) have achieved impressive visual realism and large-scale physics.
 Yet their success does not imply physical understanding: most are trained to match appearances rather than to model underlying physical state and dynamics.
@@
 Learning such models requires dense, temporally aligned signals of geometry and motion, which are extremely difficult to obtain from real-world data at scale.
-Existing 4D datasets(Wen et al., 2025a; Zhai et al., 2024) therefore tend to be appearance-rich but physics-poor, often dominated by camera motion.
+Existing 4D datasets(Wen et al., 2025; Zhai et al., 2024) therefore tend to be appearance-rich but physics-poor, often dominated by camera motion.
 To address this limitation, we leverage the diversity and controllability of physics-based simulation to obtain large amounts of dynamic video data with accurate ground-truth geometry and motion, providing supervision that is otherwise inaccessible.
@@
 To evaluate whether a model behaves as a physics-consistent 4D world model, we introduce a comprehensive evaluation protocol that combines content-level physics assessment and 4D world-level physics analysis.
-Specifically, in addition to established video-based physics benchmarks such as PhysicsIQ(Motamed et al., 2025b), which evaluate physical plausibility from visual and semantic cues, we propose a 4D evaluation suite that probes geometry and motion world representations through trajectory-based analysis.
+Specifically, in addition to established video-based physics benchmarks such as PhysicsIQ(Motamed et al., 2025), which evaluate physical plausibility from visual and semantic cues, we propose a 4D evaluation suite that probes geometry and motion world representations through trajectory-based analysis.
 Together, these complementary evaluations assess physical plausibility from what a model shows to how the underlying world evolves, providing a holistic diagnosis of physics-consistent 4D modeling beyond appearance and reconstruction quality.
@@
 Recent video generation models(OpenAI, 2024; Google DeepMind, 2024) achieve strong visual realism and temporal coherence, but are largely optimized for appearance rather than physical world modeling.
-Several works attempt to introduce physical cues into video generation, such as modeling object dynamics(Zhang et al., 2024; Aira et al., 2024) or action-conditioned motion(Chang et al., 2024; Yuan et al., 2025), and a number of benchmarks(Motamed et al., 2025b; Meng et al., 2025; Bansal et al., 2024b) evaluate physical plausibility based on visual or semantic signals.
+Several works attempt to introduce physical cues into video generation, such as modeling object dynamics(Zhang et al., 2024; Aira et al., 2024) or action-conditioned motion(Wang et al., 2025; Yuan et al., 2025), and a number of benchmarks(Motamed et al., 2025; Meng et al., 2025; Bansal et al., 2024) evaluate physical plausibility based on visual or semantic signals.
 While these approaches can improve or assess appearance-level physical realism, they operate primarily in image space and do not evaluate whether a model maintains coherent geometry, motion, or a consistent physical state over time.
@@
 Physics-Aware 4D Modeling
-A growing body of work explores incorporating physical priors into generative models, including post hoc physical simulation after generation (Stomakhin et al., 2013b; Chen et al., 2025a; Xie et al., 2024), simulation-guided generation with predefined dynamics (Yuan et al., 2023; Liu and others, 2024a; Li et al., 2025), and learning physical priors from data (Li et al., 2024; Xu et al., 2024; Yang et al., 2025a).
+A growing body of work explores incorporating physical priors into generative models, including post hoc physical simulation after generation (Stomakhin et al., 2013; Chen et al., 2025a; Xie et al., 2024), simulation-guided generation with predefined dynamics (Yuan et al., 2023; Liu et al., 2024; Li et al., 2025), and learning physical priors from data (Li et al., 2024; Xu et al., 2024; Yang et al., 2025a).
 These methods typically focus on object-level or short-horizon dynamics, rely on specialized 3D representations, or remain optimization-based, limiting scalability and generalization.
@@
 We present a physics-based simulation pipeline designed to address the limited physical fidelity and diversity of existing training data for 4D world modeling.
-In contrast to prior datasets that focus on narrow physical regimes or specific simulation paradigms (Huang et al., 2025a; Chang et al., 2024), Phys4D supports large-scale data generation with diverse object types, rich physical interactions, and accurate geometry and motion supervision.
+In contrast to prior datasets that focus on narrow physical regimes or specific simulation paradigms (Huang et al., 2025a; Wang et al., 2025), Phys4D supports large-scale data generation with diverse object types, rich physical interactions, and accurate geometry and motion supervision.
 Figure 2 provides an overview of the simulation pipeline and the range of supported object categories.
@@
 Benchmark.
-We evaluate on Physics-IQ (Motamed et al., 2025a), a benchmark designed to assess physical understanding in generative video models.
+We evaluate on Physics-IQ (Motamed et al., 2025), a benchmark designed to assess physical understanding in generative video models.
 The benchmark comprises 198 test scenarios covering diverse physical phenomena.
@@
 Implementation.
-We apply Phys4D to three backbone: WAN2.2-5b (Wan et al., 2025), CogVideoX-5b (Yang et al., 2025b), and Open-Sora V1.2 (Zheng et al., 2024).
+We apply Phys4D to three backbone: WAN2.2-5b (Wan et al., 2025), CogVideoX-5b (Yang et al., 2025b), and Open-Sora V1.2 (Zheng et al., 2024b).
 For evaluation, we generate 5-second videos conditioned on the switch frame following the Physics-IQ protocol.
@@
 Notably, flow matching is not restricted to Gaussian endpoints or linear paths and can be generalized to interpolate between arbitrary distributions (Albergo and Vanden-Eijnden, 2023; Liu et al., 2023).
-This flexibility has recently enabled efficient large-scale generative modeling, including applications to high-dimensional and structured data such as videos, where flow-based formulations offer favorable trade-offs between sampling efficiency and temporal coherence (Xu et al., 2022; Lin and others, 2024b; Peng and others, 2025).
+This flexibility has recently enabled efficient large-scale generative modeling, including applications to high-dimensional and structured data such as videos, where flow-based formulations offer favorable trade-offs between sampling efficiency and temporal coherence (Xu et al., 2022; Lin and others, 2024; Zheng et al., 2024a).
 F.2 Denoising as a Markov Decision Process
@@
 where $\hat{A}_{t}$ denotes the estimated advantage.
-This objective stabilizes policy updates and has been shown effective for fine-tuning diffusion- and flow-based generative models (Black and others, 2023; Ren and others, 2024; Liu and others, 2025).
+This objective stabilizes policy updates and has been shown effective for fine-tuning diffusion- and flow-based generative models (Black and others, 2023; Ren and others, 2024; Liu et al., 2025).
 Appendix G Related Work
@@
 Recent large-scale video generation models, such as Sora (OpenAI, 2024), Veo (Google DeepMind, 2024), and MovieDiffusion (Polyak et al., 2024), have achieved impressive visual realism and long-horizon coherence, motivating their use as foundations for generative world modeling.
-Recent world model systems further extend these capabilities toward persistent and interactive environments, either by explicitly constructing 3D scenes or by augmenting video generation with memory and feedback mechanisms (Chen et al., 2025d; Agarwal and others, 2025; Yu et al., 2024a; Wang and others, 2025; Team and others, 2025).
+Recent world model systems further extend these capabilities toward persistent and interactive environments, either by explicitly constructing 3D scenes or by augmenting video generation with memory and feedback mechanisms (Chen et al., 2025c; Agarwal and others, 2025; Yu et al., 2024; Huang et al., 2025b; Team and others, 2025).
 While these approaches emphasize scalability, global coherence, and interactivity, the underlying physical state is often implicit, making fine-grained physical consistency—such as local geometry–motion coupling and precise object interactions—difficult to enforce.
@@
 Recent works further scale up 4D datasets by exploiting in-the-wild videos and automatic annotation pipelines.
-DynamicVerse introduces a large-scale multimodal 4D dataset derived from internet videos, providing metric-scale geometry, camera parameters, instance-level masks, motion cues, and hierarchical textual descriptions across over 100K dynamic scenes (Wen et al., 2025b).
+DynamicVerse introduces a large-scale multimodal 4D dataset derived from internet videos, providing metric-scale geometry, camera parameters, instance-level masks, motion cues, and hierarchical textual descriptions across over 100K dynamic scenes (Wen et al., 2025).
 Complementary datasets such as OmniWorld and Ego-Exo4D extend this direction by covering diverse environments, viewpoints, and sensor modalities, facilitating evaluation of general-purpose 4D world understanding models (Zhou et al., 2025; Grauman et al., 2023).
@@
 To facilitate standardized evaluation, several benchmarks have been proposed to assess general video generation quality.
-VBench (Huang and others, 2024) and EvalCrafter (Liu and others, 2024b) provide comprehensive evaluation frameworks covering dimensions such as visual fidelity, diversity, temporal coherence, and scalability.
+VBench (Huang and others, 2024) and EvalCrafter (Liu and others, 2024) provide comprehensive evaluation frameworks covering dimensions such as visual fidelity, diversity, temporal coherence, and scalability.
 T2V-CompBench (Sun and others, 2024) further focuses on compositionality, testing whether generated videos faithfully follow complex and multi-object textual prompts.
@@
 Recent studies highlight the importance of physical realism in video generation and propose dedicated physics-oriented benchmarks.
-VideoPhy (Bansal et al., 2024a) evaluates physical commonsense reasoning in generated videos, while PhyGenBench (Meng and others, 2024) assesses adherence to basic physical laws such as object permanence and gravity.
-DEVIL (Liao and others, 2024) emphasizes dynamic realism, focusing on motion continuity and temporal consistency, and Physics-IQ (Motamed et al., 2025b) further probes understanding of fundamental physical principles including fluid dynamics and thermodynamics.
+VideoPhy (Bansal et al., 2024) evaluates physical commonsense reasoning in generated videos, while PhyGenBench (Meng and others, 2024) assesses adherence to basic physical laws such as object permanence and gravity.
+DEVIL (Liao and others, 2024) emphasizes dynamic realism, focusing on motion continuity and temporal consistency, and Physics-IQ (Motamed et al., 2025) further probes understanding of fundamental physical principles including fluid dynamics and thermodynamics.
 Despite these advances, existing benchmarks are often limited in scenario diversity or focus on a narrow set of physical laws.
-Motivated by these limitations, recent efforts introduce substantially expanded benchmarks that evaluate video generation models across a broader range of physical phenomena, including both realistic and deliberately unphysical dynamics (Meng and others, 2025).
+Motivated by these limitations, recent efforts introduce substantially expanded benchmarks that evaluate video generation models across a broader range of physical phenomena, including both realistic and deliberately unphysical dynamics.
 Empirical results show that state-of-the-art models perform significantly worse under identical settings, revealing the increased difficulty and diagnostic value of physics-aware benchmarks and underscoring the need for models that reason about physical consistency beyond visual realism.
@@
 This line of work first generates static images or 3D assets using generative models, and subsequently applies external physics simulators to animate them into dynamic scenes or videos.
-Representative approaches employ particle- or continuum-based simulators, such as the Material Point Method (MPM), to model object deformation and motion (Stomakhin et al., 2013a).
-Recent studies adopt this paradigm to produce physically plausible dynamics by post-processing generated content with physics engines (Lin and others, 2024a; Hsu and others, 2024).
+Representative approaches employ particle- or continuum-based simulators, such as the Material Point Method (MPM), to model object deformation and motion (Stomakhin et al., 2013).
+Recent studies adopt this paradigm to produce physically plausible dynamics by post-processing generated content with physics engines.
 While this approach offers explicit and controllable physical behavior, it typically requires substantial manual effort and task-specific simulator design, limiting scalability and generalization.
@@
 Another category of methods first performs physical simulation to obtain plausible motion trajectories or object states, which are then used as conditional inputs for video generation models.
-For example, PhysGen (Lin and others, 2024a) segments dynamic objects from an input image, simulates their motion using Newtonian mechanics, and conditions a video generator on both simulated trajectories and static backgrounds.
+For example, PhysGen (Liu et al., 2024) segments dynamic objects from an input image, simulates their motion using Newtonian mechanics, and conditions a video generator on both simulated trajectories and static backgrounds.
 Although this strategy improves controllability and motion realism, the generative model itself does not possess intrinsic physical reasoning capability, and physical rules and parameters must be manually specified for each scenario.
 Generation with Learned Physics Priors.
-More recent approaches aim to incorporate physics priors directly into the generative process by leveraging learned or pretrained models (Chang et al., 2024; Zhang and others, 2025; Chefer and others, 2025).
+More recent approaches aim to incorporate physics priors directly into the generative process by leveraging learned or pretrained models (Wang et al., 2025; Zhang and others, 2025).
 These methods encode physical knowledge as latent dynamics, auxiliary supervision, or physics-aware evaluators.